### v1

# HiCache: Training-free Acceleration of Diffusion Models via Hermite Polynomial-based Feature Caching

Abstract
Diffusion models have achieved remarkable success in content generation but suffer from prohibitive computational costs due to iterative sampling. While recent feature caching methods tend to accelerate inference through temporal extrapolation, these methods still suffer from server quality loss due to the failure in modeling the complex dynamics of feature evolution. To solve this problem, this paper presents HiCache, a training-free acceleration framework that fundamentally improves feature prediction by aligning mathematical tools with empirical properties. Our key insight is that feature derivative approximations in Diffusion Transformers exhibit multivariate Gaussian characteristics, motivating the use of Hermite polynomials—the potentially theoretically optimal basis for Gaussian-correlated processes. Besides, We further introduce a dual-scaling mechanism that ensures numerical stability while preserving predictive accuracy. Extensive experiments demonstrate HiCache’s superiority: achieving 6.24× speedup on FLUX.1-dev while exceeding baseline quality, maintaining strong performance across text-to-image, video generation, and super-resolution tasks. Core implementation is provided in the appendix, with complete code to be released upon acceptance.
Figure 1: Visual demonstration of HiCache using the FLUX.1-dev model, with a $\boldsymbol{6.24\times}$ speedup ($N_{\text{interval}}=7$).
1. Introduction
Diffusion Models (DMs) (Ho, Jain, and Abbeel 2020; Nichol and Dhariwal 2021; Dhariwal and Nichol 2021), especially when implemented with the scalable Transformer architecture (DiT) (Peebles and Xie 2023), have set a new standard for high-fidelity content generation (Chen et al. 2023; Rombach et al. 2022; Karras et al. 2022; Saharia et al. 2022; Balaji et al. 2022). This success, however, is tempered by the substantial computational overhead inherent to the diffusion process. The iterative sampling, requiring hundreds of sequential forward passes through large models, results in high latency that critically hinders practical deployment. Consequently, developing efficient acceleration methods is a key research objective.
Among various acceleration strategies, feature caching has emerged as a particularly effective and training-free approach. Motivated by the observation that diffusion models exhibit similar features in the adjacent timesteps, the original feature caching seeks to directly reuse features in the adjacent timesteps, which leads to significant generation quality loss when the acceleration ratios increase (Selvaraju et al. 2024; Ma et al. 2023; Zou et al. 2024). Recently, the paradigm has advanced to a “cache-then-forecast” mechanism as TaylorSeer (Liu et al. 2025), where Taylor series expansion is used to extrapolate features in the future timesteps, significantly reducing the error from feature caching. TaylorSeer significantly reduces the error introduced by feature reusing, but still faces fundamental theoretical limitations: The standard polynomial basis of a Taylor series is suboptimal for modeling the complex, non-monotonic trajectories of feature evolution in DMs.
Figure 2 visualize the features of the original diffusion models (i.e., ground-truth) and the predictive features from TaylorSeer. It is observed that Taylor-based extrapolation exhibits significant deviations from ground truth trajectories, particularly at turning points where its monotonic nature fails to capture the underlying dynamics. This mismatch between the Taylor-based predictive model and the underlying feature distribution establishes a bottleneck for achieving much higher acceleration ratios without significant quality degradation. Besides, TaylorSeer is designed to employ high-order Taylor expansion to make use of more historical information, making it sensitive to the noise in the feature trajectory, further limiting its effectiveness.
Figure 2: Trajectory prediction comparison of Taylor and HiCache methods on FLUX. “Full trajectory” indicates the feature trajectory from the original Flux (Esser et al. 2024) in the 14th and 28th layer. The y-axis denotes the most principal component of the features in diffusion models. HiCache demonstrates more stable fitting at trajectory turning points, while Taylor expansion tends to over-extrapolate and deviate from original trajectories.
To overcome this limitation, in this paper, we propose HiCache (Hermite Polynomial-based Feature Cache), a framework that improves the feature forecasting by aligning the predictive basis with the intrinsic statistical properties of the feature dynamics. Our approach is based on a key empirical observation: the derivative approximations of feature vectors in DiTs consistently exhibit a multivariate Gaussian distribution (details in Section 3.3.1). Motivated by this finding, HiCache replaces the standard polynomial basis with scaled Hermite polynomials. According to approximation theory (Szegö 1975), Hermite polynomials are the potentially optimal orthogonal basis for representing Gaussian-correlated processes, making them theoretically better suited for this task. Besides, to address the numerical challenges of Hermite polynomials at large extrapolation steps, we further introduce a dual-scaling mechanism that simultaneously constrains predictions within the stable oscillatory regime and suppresses exponential coefficient growth in high-order terms through a single hyperparameter. This principled design enables both accurate and numerically stable feature prediction. As shown in Figure 2, the feature trajectory predicted by HiCache closely matches the original FLUX, demonstrating its great effectiveness.
[…]
The main contributions of this work are as follows:
  • We systematically validate the multivariate Gaussian nature of feature derivative approximations in Diffusion Transformers, offering a new statistical foundation for designing more efficient feature caching methods.
  • We propose HiCache, which introduces Hermite polynomials into the feature caching of diffusion models, and propose a dual-scaling mechanism to simultaneously constrain predictions within the stable oscillatory regime and suppress exponential coefficient growth in high-order terms, achieving robust numerical stability.
  • We conduct extensive experiments on four diffusion models and generative tasks, demonstrating HiCache’s universal superiority and broad applicability.
2. Related Work
Sampling Step Reduction
aims to reduce the total number of denoising iterations. DDIM (Song, Meng, and Ermon 2021) introduced deterministic sampling with larger steps, while DPM-Solver (Lu et al. 2022) formalized this as an ODE with high-order numerical methods. Recent advances include knowledge distillation (Meng et al. 2023) and Consistency Models (Song et al. 2023) for few-step generation. Other notable approaches include cascaded diffusion models (Ho et al. 2022a) and improved sampling techniques (Karras et al. 2022). Text-to-image diffusion models have seen remarkable progress (Saharia et al. 2022; Balaji et al. 2022; Ramesh et al. 2022; Podell et al. 2023; Zhang, Rao, and Agrawala 2023), while video generation (Singer et al. 2022; Ho et al. 2022b) presents additional temporal challenges.
Denoising Network Compression
focuses on reducing per-step computational cost. Model compression methods include neural network pruning (Fang et al. 2023) and quantization (Li et al. 2023) have applied the traditional model compression methods to diffusion models. Besides, token merging (Bolya and Hoffman 2023) has been introduced to reduce the number of tokens computed in each timestep. However, these methods tend to lead clear loss in generation quality, and sometimes rely on retraining the diffusion model.
Feature Caching
has been recently proposed to offer a training-free alternative. Motivated by the great similarity of features in the adjacent timesteps, these methods aim to skip the computation in some timesteps based on the historical features, which can be roughly divided into two paradigms: (1) Cache-then-Reuse directly reuses features from adjacent timesteps (DeepCache (Ma et al. 2023), FORA (Selvaraju et al. 2024), ToCa (Zou et al. 2024)), but suffers quality degradation at large intervals; (2) Cache-then-Forecast predicts future features using historical data. TaylorSeer (Liu et al. 2025) pioneered this approach with Taylor series expansion, significantly improving quality at high acceleration ratios. However, its standard polynomial basis remains suboptimal for complex feature dynamics, motivating our Hermite polynomial-based approach.
3. Methodology
3.1 Preliminaries: Diffusion Transformer
The Diffusion Transformer (DiT) architecture is central to many state-of-the-art diffusion models. Building upon the self-attention mechanism (Vaswani et al. 2017), it operates by progressively denoising a noisy input $x_{t}$ through a series of $L$ stacked blocks. Within each block, an input feature map $\mathcal{F}_{\text{in}}$ is updated via sequential residual corrections:
$\mathcal{F}^{\prime}$ $=\mathcal{F}_{\text{in}}+\text{MHSA}(\text{AdaLN}(\mathcal{F}_{\text{in}},t))$ (1)
$\mathcal{F}_{\text{out}}$ $=\mathcal{F}^{\prime}+\text{MLP}(\text{AdaLN}(\mathcal{F}^{\prime},t))$ (2)
where AdaLN denotes adaptive layer normalization conditioned on timestep $t$. This temporal conditioning, combined with nonlinearities (e.g., Softmax and GeLU), causes feature trajectories to exhibit complex, non-monotonic dynamics—a significant challenge for acceleration techniques relying on smooth extrapolation. The importance of such architectural choices in generative models has been widely studied (Karras et al. 2022; Dhariwal and Nichol 2021).
3.2 Taylor Expansion-based Feature Prediction
To accelerate sampling, existing methods like TaylorSeer extrapolate future features using Taylor series. We first formalize the finite difference framework:
Definition 1 (Finite Difference Operator).
The $i$-th order backward difference operator $\Delta^{(i)}$ on feature trajectory $\{\mathcal{F}_{t}\}$ is defined recursively:
$\Delta^{i}\mathcal{F}_{t}=\frac{\Delta^{i-1}\mathcal{F}_{t}-\Delta^{i-1}\mathcal{F}_{t-N_{\text{interval}}}}{N_{\text{interval}}},\quad\text{with}\quad\Delta^{0}\mathcal{F}_{t}=\mathcal{F}_{t}$ (3)
Taylor expansion-based prediction then takes the form:
$\hat{\mathcal{F}}_{t-k}^{\text{Taylor}}=\mathcal{F}_{t}+\sum_{i=1}^{m}\frac{\Delta^{i}\mathcal{F}_{t}}{i!}(-k)^{i}$ (4)
Proposition 1 (Limitation of Monomial Basis).
For feature trajectories with bounded variation but containing turning points, the Taylor prediction error grows as:
$\|\hat{\mathcal{F}}_{t-k}^{\text{Taylor}}-\mathcal{F}_{t-k}\|=O\left(\frac{k^{m+1}}{(m+1)!}\sup_{\xi\in[t-k,t]}\|\mathcal{F}^{(m+1)}(\xi)\|\right)$ (5)
where the supremum can be arbitrarily large at trajectory inflection points.
This limitation of the monomial basis $\{(-k)^{i}\}_{i=1}^{m}$ motivates our search for a more suitable basis function.
3.3 HiCache: Hermite-based Feature Caching
In this paper, we propose to leverage the Gaussian nature of feature dynamics for optimal basis selection.
Observation 1 (Gaussianity of Feature Derivatives).
Let $\{\mathcal{F}_{t}\}_{t=1}^{T}$ be the feature trajectory in a Diffusion Transformer. The finite difference approximations $\{\Delta^{k}\mathcal{F}_{t}\}$ follow a multivariate Gaussian distribution:
$\Delta^{k}\mathcal{F}_{t}\sim\mathcal{N}(\mu_{k},\Sigma_{k})$ (6)
Validated via Energy statistical tests with p-value = 1.0 across all configurations (see Section 5.1).
Corollary 1 (Optimal Basis Selection).
By the Karhunen-Loève theorem, Hermite polynomials $\{H_{n}\}$ form the optimal orthogonal basis for Gaussian-correlated processes (Narcowich, Ward, and Wendland 2005), implying superior approximation properties over monomials.
3.3.2 Hermite Polynomials as Basis Functions
Given the Gaussian characteristics, we adopt Hermite polynomials as our basis functions. The physicist’s Hermite polynomials $H_{n}(x)$ (Abramowitz and Stegun 1964) are defined via the Rodrigues formula:
$H_{n}(x)=(-1)^{n}e^{x^{2}}\frac{d^{n}}{dx^{n}}e^{-x^{2}}$ (7)
and satisfy the recurrence relation:
$H_{n+1}(x)=2xH_{n}(x)-2nH_{n-1}(x)$ (8)
Unlike monotonically growing Taylor basis functions, Hermite polynomials exhibit inherent oscillatory behavior (as shown in Figure 3). This oscillatory nature provides implicit regularization (van der Westhuizen and Lasenby 2018), enabling more accurate capture of non-monotonic dynamics in neural network optimization. However, Hermite polynomials face numerical challenges: they grow exponentially for large $|x|$ and have rapidly increasing coefficients due to the factor of 2 in the recurrence. To address this, we introduce scaled Hermite polynomials:
Definition 2 (Scaled Hermite Basis).
Scaled Hermite polynomials with contraction factor $\sigma\in(0,1)$ are defined as:
$\tilde{H}_{n}(x)=\sigma^{n}H_{n}(\sigma x)$ (9)
This scaling provides dual stabilization: (i) input scaling $\sigma x$ constrains predictions within the stable oscillatory regime, and (ii) coefficient scaling $\sigma^{n}$ suppresses exponential growth in high-order terms.
Figure 3: Systematic comparison of Taylor and Hermite basis functions Hermite’s oscillatory behavior (e.g., $H_{2}(x)$ negative offset) captures non-monotonic evolution better than monotonic Taylor growth. (Y-axis: symmetric log scale).
3.3.3 HiCache Algorithm
Proposition 2 (HiCache Feature Prediction).
Given cached derivative approximations $\{\Delta^{i}\mathcal{F}_{t}\}_{i=0}^{N_{\text{order}}}$, the optimal predictor under Gaussian feature dynamics is:
$\hat{\mathcal{F}}_{t-k}^{\text{HiCache}}=\mathcal{F}_{t}+\sum_{i=1}^{N_{\text{order}}}\frac{\Delta^{i}\mathcal{F}_{t}}{i!}\tilde{H}_{i}(-k)$ (10)
where $k\in\{1,\ldots,N_{\text{interval}}-1\}$ is the prediction horizon and $\tilde{H}_{i}$ is the scaled Hermite polynomial from Definition 2.
Proof Intuition. Under the empirically validated Gaussian assumption (Observation 1), Corollary 1 suggests that Hermite polynomials provide a theoretically motivated basis for feature prediction. The factorial normalization ensures proper scaling of higher-order terms. See Appendix A.6 for detailed theoretical analysis and supporting arguments.
Algorithm 1 HiCache: Hermite-based Feature Caching
0: Interval $N_{\text{interval}}$, order $N_{\text{order}}$, contraction $\sigma$
1:for timestep $t=T$ to 11 do
2:  if $t\mod N_{\text{interval}}=0$ then
3:   Compute $\mathcal{F}_{t}=\Phi_{t}(x_{t})$ {Full model forward}
4:   Update cache: $\{\Delta^{i}\mathcal{F}_{t}\}_{i=0}^{N_{\text{order}}}$ via Def. 1
5:  else
6:   $k\leftarrow t\mod N_{\text{interval}}$
7:   Predict $\hat{\mathcal{F}}_{t}$ using Eq. (10)
8:  end if
9:end for
Computational Complexity. HiCache reduces computational cost by factor $(N_{\text{interval}}-1)/N_{\text{interval}}$ while maintaining $O(N_{\text{order}})$ prediction overhead per skipped step.
Proposition 3 (Error Bounds for HiCache).
The total prediction error of HiCache decomposes as:
$\mathbf{E}_{\text{total}}=\mathbf{E}_{\text{truncation}}+\mathbf{E}_{\text{approximation}}+\mathbf{E}_{\text{numerical}}$ (11)
with the bound:
$\|\mathbf{E}_{\text{total}}\|\leq\underbrace{O\left(\frac{(\sigma\sqrt{2}|\Delta s|)^{N+1}}{\sqrt{(N+1)!}}\right)}_{\text{Hermite truncation}}+\underbrace{O(\Delta t_{\text{hist}}\sqrt{N})}_{\text{Finite difference}}+\underbrace{O(\epsilon_{\text{machine}})}_{\text{Floating point}}$ (12)
Corollary 2 (Superiority over Taylor Expansion).
HiCache achieves:
  1. 1.
    Faster convergence: $O(1/\sqrt{(N+1)!})$ vs. $O(1/(N+1)!)$
  2. 2.
    Exponential suppression via $\sigma^{N+1}$
  3. 3.
    Automatic adaptation to Gaussian feature statistics
Proof. See Appendix A.3 for complete derivation.
Figure 4: An overview of HiCache compared with TaylorSeer. TaylorSeer (orange) predicts the feature with a power basis. HiCache (green) uses the same functional form but replaces the basis with Hermite polynomials. Dashed curves show predicted trajectories against ground truth, highlighting error accumulation at skipped steps.
4. Experiments
4.1 Experiment Settings
We conduct comprehensive experiments across four representative generative tasks using state-of-the-art diffusion models: text-to-image generation with FLUX.1-dev (Esser et al. 2024), text-to-video generation with HunyuanVideo (Kong et al. 2024), class-conditional image generation with DiT-XL/2 (Peebles and Xie 2023), and image super-resolution with a modified Inf-DiT (Yang et al. 2024). Each task employs standard evaluation protocols and widely-adopted benchmarks to ensure fair comparison with existing methods. More detailed information is provided in Appendix A.4.
4.2 Text-to-Image Generation
Quantitative Analysis.
As shown in Table 1, HiCache demonstrates a clear quantitative advantage. We evaluate using ImageReward (Xu et al. 2023) as the primary metric for perceptual quality assessment. At a  5.55x acceleration, HiCache not only surpasses its main competitor, TaylorSeer, but also slightly exceeds the unaccelerated baseline, suggesting near-lossless performance. This advantage becomes more pronounced at higher accelerations ( 6.24x), where HiCache maintains a robust score while competitors’ performance degrades more significantly.
Table 1: Quantitative comparison on the text-to-image generation task with the FLUX.1-dev model.
Method Latency(s) $\downarrow$ Speed $\uparrow$ FLOPs(T) $\downarrow$ Speed $\uparrow$ Image Reward $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CLIP $\uparrow$
[…]
Numbers in parentheses show the deviation from the DDIM-50 baseline (Image Reward=0.9872). PSNR and LPIPS are compared against the ideal values of $\infty$ and 0, respectively. Among the compared methods, only ToCa does not utilize efficient attention.
Qualitative Analysis.
As illustrated in Figure 5, HiCache consistently produces high-fidelity results, avoiding common artifacts like facial distortions, conceptual errors, or text inaccuracies that affect competing methods. Furthermore, Figure 6 highlights HiCache’s superior ability to preserve high-frequency details, such as intricate textures and subtle reflections, which are often lost in other accelerated outputs.
Figure 5: Qualitative comparison on the text-to-image task across nine diverse prompts.
Figure 6: Visualization of high-frequency detail preservation.
Robustness Under Increasing Acceleration.
Figure 7 assesses stability under increasing acceleration factors (from 1.00x to 9.00x). TaylorSeer’s output shows significant style drift and structural degradation, consistent with the error accumulation discussed in Sec. 5.1. In contrast, HiCache maintains high stability and aesthetic consistency, demonstrating its robustness against such errors even at the highest acceleration levels.
Figure 7: Comparison between TaylorSeer (second-best method) and HiCache in different acceleration ratios.
4.3 Text-to-Video Generation
Quantitative Analysis.
We further evaluate HiCache on text-to-video generation using the HunyuanVideo model (Kong et al. 2024). As detailed in Table 3, which benchmarks performance with VBench (Huang et al. 2024), HiCache consistently outperforms competitors. Notably, when comparing higher-order configurations (e.g., $N=7,O=2$), HiCache demonstrates even more significant advantages over TaylorSeer, which aligns with our theoretical analysis in Section 3.3.2 that Hermite polynomials’ oscillatory behavior better captures complex feature dynamics at higher orders, while Taylor expansion’s monotonic nature leads to increasing over-extrapolation.
Table 2: Quantitative comparison on class-to-image generation on ImageNet with DiT-XL/2.
Method Efficient Attention Latency(s) $\downarrow$ FLOPs(T) $\downarrow$ Speed $\uparrow$ FID $\downarrow$ sFID $\downarrow$ Inception Score $\uparrow$
[…]
Numbers in parentheses show the deviation from the DDIM-50 baseline (FID=2.32, sFID=4.32, IS=241.25).
Table 3: Quantitative comparison on text-to-video generation with HunyuanVideo on Vbench.
Method Lat.(s) $\downarrow$ Spd. $\uparrow$ FLOPs(T) $\downarrow$ Spd. $\uparrow$ VBench(%) $\uparrow$
[…]
Note: All compared methods utilize efficient attention.
Qualitative Analysis. The qualitative results in Figure 8 reveal that competing methods often suffer from temporal inconsistencies, leading to artifacts such as distorted objects and incoherent movements, a common challenge in video generation (Ho et al. 2022b; Singer et al. 2022). In contrast, HiCache maintains high temporal stability and logical consistency, producing more realistic and coherent video sequences.
[…]
4.4 Class-Conditional Image Generation
Quantitative Analysis.
On the task of class-conditional ImageNet generation (Dhariwal and Nichol 2021) with the DiT-XL/2 model (Peebles and Xie 2023), HiCache again demonstrates superior performance. As shown in Table 2, it achieves a lower (better) FID score (Salimans et al. 2016) compared to competing methods at  5x acceleration. This performance gap widens at higher acceleration ratios ( 7.1x), where HiCache maintains a strong FID score while others degrade significantly, further confirming its robustness in high-acceleration scenarios.
4.5 Super-Resolution
On the image super-resolution task with a patch-based Inf-DiT model (Yang et al. 2024), HiCache achieves a significant theoretical ( 5.93×) and wall-clock ( 2.43×) speedup while maintaining near-lossless quality. This task follows the NTIRE 2025 challenge protocol (Chen et al. 2025). As shown in Figure 9, both PSNR and SSIM consistently outperform TaylorSeer, with particularly notable improvements in SSIM. This demonstrates superior stability and quality preservation at high acceleration (detailed results in Appendix A.5).
Figure 9: Performance comparison of HiCache and TaylorSeer on image super-resolution across different acceleration intervals. HiCache consistently outperforms TaylorSeer in PSNR and SSIM, demonstrating superior stability and quality preservation at higher acceleration ratios.
4.6 Ablation Study
To validate the necessity of the contraction factor $\sigma$ introduced in Section 3.3.2, we conduct an ablation study using the same experimental settings as Section 4.2. As shown in Table 4, the results confirm that the contraction mechanism is essential for the method’s viability, with the optimal range representing a careful balance between maintaining the oscillatory properties of Hermite basis functions and ensuring numerical stability, validating our theoretical design that addresses both prediction domain constraints and exponential coefficient growth suppression.
Table 4: Ablation study demonstrating the necessity of the contraction factor $\sigma$ on FLUX for text-to-image generation.
Method FLOPs Speed $\uparrow$ Image Reward $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
[…]
Note: $\sigma=1.0$ means no contraction factor is applied. Without contraction, HiCache performs worse than TaylorSeer, proving that the dual-scaling mechanism is essential for method viability.
5. Discussion
Empirical Validation of Observation 1
The foundation of HiCache rests on Observation 1, which claims that feature derivative approximations follow a multivariate Gaussian distribution. We validated this using Energy tests (Székely and Rizzo 2005) across 1st to 5th order feature differences in five key FLUX model (Esser et al. 2024) modules. The tests confirmed Gaussianity with maximum confidence (p-value = 1.0) across all 25 configurations, providing strong empirical support for our theoretical framework and justifying Corollary 1 that Hermite polynomials form the optimal basis.
Prediction Simulation Framework
To fairly evaluate HiCache against Taylor expansion, we established a simulation framework using real feature trajectories from the FLUX model with activation interval $N_{\text{interval}}=6$, where derivative approximations are computed using sparsely sampled data from previous activation steps. Our simulation eliminates cumulative error by initializing each prediction task with ground-truth historical data. This design isolates the analysis from cascading inaccuracies and enables meaningful quantitative comparison of the basis functions’ intrinsic approximation capabilities, ensuring direct evaluation of performance differences between Hermite and Taylor bases.
Empirical Verification of Oscillatory Advantages
Figure 10 provides empirical evidence for the theoretical advantages of Hermite polynomials’ oscillatory properties (discussed in Section 3.3.2). Using the non-cumulative error framework on representative FLUX features, we observe that while both methods perform similarly at order 1, Taylor expansion increasingly over-extrapolates at higher orders, especially at trajectory turning points—precisely confirming Proposition 1. HiCache’s oscillatory Hermite basis, stabilized by the contraction factor $\sigma$ (Definition 2), maintains accurate fitting across all orders.
Figure 10: Visualization example of trajectory prediction for different polynomial orders under FLUX architecture. HiCache demonstrates more stable fitting at trajectory turning points, while Taylor expansion over-extrapolates.
Quantitative Validation of Proposition 3
To empirically verify the tighter error bounds promised by Proposition 3, we quantitatively compared HiCache against Taylor expansion using the simulation framework. We define the relative error ratio as $R=\nicefrac{{\text{Taylor MSE}}}{{\text{Hermite MSE}}}$ (where $R>1$ indicates superior HiCache performance) and analyzed 25 configurations (5 FLUX modules × 5 polynomial orders).
Figure 11 presents the cumulative error ratios:
Figure 11: Cumulative error ratios (Taylor MSE / HiCache MSE) across 5 FLUX modules as prediction step size varies. HiCache consistently outperforms TaylorSeer (all ratios >1.0>1.0), with the advantage becoming more pronounced at larger prediction steps and higher polynomial orders.
The results empirically confirm HiCache’s consistent advantage ($R>1.0$ across all modules), with the performance gap widening for larger prediction steps and higher polynomial orders. This validates our theoretical analysis: the $O(1/\sqrt{(N+1)!})$ convergence rate from Proposition 3 indeed provides tighter bounds than Taylor’s $O(1/(N+1)!)$, with the advantage becoming more pronounced as $N$ increases—precisely as predicted by Corollary 2.
6. Conclusion
We presented HiCache, a training-free acceleration framework for diffusion models that fundamentally improves feature prediction by replacing Taylor series with scaled Hermite polynomials. Our approach is grounded in rigorous theory: Observation 1 establishes the Gaussian nature of feature derivatives, Corollary 1 identifies Hermite polynomials as the optimal basis, and Proposition 3 provides theoretical guarantees of superior convergence. The scaled Hermite basis (Definition 2) with dual-scaling mechanism $\tilde{H}_{n}(x)=\sigma^{n}H_{n}(\sigma x)$ ensures numerical stability through a single hyperparameter, enabling HiCache to achieve superior performance across diverse tasks:  6.24× speedup on FLUX.1-dev while exceeding baseline quality, maintaining strong FID scores at high acceleration ratios, and pioneering cache-based acceleration for super-resolution. As a plug-and-play solution requiring no model retraining, HiCache establishes a new paradigm for diffusion acceleration by demonstrating that matching basis functions to intrinsic data statistics yields both theoretical and practical benefits.
A.1 Hermite Polynomials: Mathematical Foundations
We adopt the physicist’s Hermite polynomials $H_{n}(x)$, whose standard differential form (Rodrigues formula) is:
$H_{n}(x)=(-1)^{n}e^{x^{2}}\frac{d^{n}}{dx^{n}}e^{-x^{2}}$ (13)
The Gaussian kernel $e^{-x^{2}}$ in this formula directly embodies the connection with Gaussian measure. From this differential form, we can derive the recurrence relation:
$H_{n+1}(x)=2xH_{n}(x)-2nH_{n-1}(x)$ (14)
where $H_{0}(x)=1$ and $H_{1}(x)=2x$. The first few polynomials are:
$H_{0}(x)$ $=1$
$H_{1}(x)$ $=2x$
$H_{2}(x)$ $=4x^{2}-2$ (15)
$H_{3}(x)$ $=8x^{3}-12x$
$H_{4}(x)$ $=16x^{4}-48x^{2}+12$
A.2 Detailed Algorithm Description
HiCache algorithm operates through two main phases. Central to this is the concept of the derivative approximation, $\Delta^{k}\mathcal{F}$, which approximates the $k$-th order derivative of a feature $\mathcal{F}$ at a given timestep. As defined in Eq. (3), it is computed recursively from lower-order approximations at previous activation steps.
Phase 1: Cache Update (at activation steps $t_{i}$ where $t_{i}\mod N_{\text{interval}}=0$)
This phase uses finite differences to compute high-order derivative approximations and updates the cache. The interval between adjacent activation steps is $\Delta t_{\text{hist}}=N_{\text{interval}}$.
Algorithm 2 Cache Update with Derivative Approximations
0: Current feature $\mathcal{F}(t_{i})$, previous activation step $t_{i-1}$, previous cache $\{\Delta^{k}\mathcal{F}(t_{i-1})\}_{k=0}^{N_{order}}$, interval $N_{\text{interval}}$
0: Updated cache $\{\Delta^{k}\mathcal{F}(t_{i})\}_{k=0}^{N_{order}}$
[…]
3:for $k=0$ to $N_{order}-1$ do
4:  {Compute $(k+1)$-th order derivative approximation}
5:  $\Delta^{k+1}\mathcal{F}(t_{i})\leftarrow(NewCache[k]-\Delta^{k}\mathcal{F}(t_{i-1}))/\Delta t_{hist}$
6:  $NewCache[k+1]\leftarrow\Delta^{k+1}\mathcal{F}(t_{i})$
7:end for
8:return  NewCacheNewCache
Phase 2: Feature Prediction (at non-activation steps $t$ where $t\mod N_{\text{interval}}\neq 0$)
This phase performs efficient prediction based on cached derivative approximations and scaled Hermite polynomials. The prediction step $\Delta t$ ranges from $1,N_{\text{interval}}-1]$.
Algorithm 3 Feature Prediction
0: Current step $t$, most recent activation step $t_{last}$, cache $\{\Delta^{k}\mathcal{F}(t_{last})\}_{k=0}^{N_{order}}$, contraction factor $\sigma$
0: Predicted feature $\hat{\mathcal{F}}$
[…]
3:for $k=1$ to $N_{order}$ do
4:  {Compute scaled Hermite polynomial value}
5:  $h_{k}\leftarrow\tilde{H}_{k}(\Delta t)$
6:  {This involves computing $H_{k}(\sigma\cdot\Delta t)$ and scaling by $\sigma^{k}$}
7:  $\hat{\mathcal{F}}\leftarrow\hat{\mathcal{F}}+\frac{h_{k}}{k!}\cdot\Delta^{k}\mathcal{F}(t_{last})$ {Accumulate}
8:end for
9:return  $\hat{\mathcal{F}}$
where Hermite polynomials are efficiently computed through the recurrence relation:
  • $H_{0}(x)=1$
  • $H_{1}(x)=2x$
  • $H_{k+1}(x)=2xH_{k}(x)-2kH_{k-1}(x)$ for $k\geq 1$
This algorithm design balances computational efficiency with prediction accuracy. Key parameters include:
  • Interval parameter $N_{\text{interval}}\in\mathbb{N}^{+}$: Controls activation frequency, with larger values providing higher acceleration ratio but potentially increased prediction error
  • Contraction factor $\sigma\in(0,1)$: Stabilization parameter that ensures numerical stability of Hermite polynomials within the prediction interval
  • Maximum order $N_{\text{order}}\in\mathbb{N}^{+}$: Expansion order that balances approximation accuracy and computational complexity, with practical implementations typically using $N_{\text{order}}\leq 4$
A.3 Complete Error Analysis and Theoretical Guarantees
This section provides rigorous proofs for the error bounds stated in Proposition 1 (Error Bounds for HiCache) in the main paper. We assume the feature evolution trajectory $f:[0,T]\to\mathcal{F}$ is sufficiently smooth with bounded derivatives up to order $N_{\text{order}}+1$.
The total prediction error decomposes into three components:
$\mathbf{E}_{\text{total}}=\mathbf{E}_{\text{truncation}}+\mathbf{E}_{\text{approximation}}+\mathbf{E}_{\text{numerical}}$ (16)
A.3.1 Truncation Error Analysis
Based on the Karhunen-Loève theorem and validated Gaussian properties, the feature function $f(s)$ can be expanded under the weight function $w(x)=e^{-(\sigma x)^{2}}$ as:
$f(s_{\text{last}}+\Delta s)=\sum_{k=0}^{\infty}a_{k}\frac{\tilde{H}_{k}(\Delta s)}{\sqrt{2^{k}k!\sqrt{\pi}}}$ (17)
where the coefficients $a_{k}=\langle f,\tilde{H}_{k}\rangle_{w}$ possess rapid decay properties.
Lemma 1 (Truncation Error Bound).
The truncation error from using order $N_{\text{order}}$ satisfies:
$\mathbf{E}_{\text{truncation}}\leq C_{2}\frac{(\sigma\sqrt{2}|\Delta s|)^{N_{\text{order}}+1}}{\sqrt{(N_{\text{order}}+1)!}}\exp\left(\frac{(\sigma\Delta s)^{2}}{2}\right)$ (18)
Proof.
Using the asymptotic bound $|H_{k}(x)|\leq C_{1}\sqrt{2^{k}k!}e^{x^{2}/2}$ and the scaling properties of $\tilde{H}_{k}$:
$\mathbf{E}_{\text{truncation}}$ $=\left|\sum_{k=N_{\text{order}}+1}^{\infty}\frac{1}{k!}\tilde{H}_{k}(\Delta s)\,\Delta^{k}f(s_{\text{last}})\right|$ (19)
$\leq\sum_{k=N_{\text{order}}+1}^{\infty}\frac{\sigma^{k}|H_{k}(\sigma\Delta s)|}{k!}\cdot|\Delta^{k}f|$ (20)
$\leq C_{2}\frac{(\sigma\sqrt{2}|\Delta s|)^{N_{\text{order}}+1}}{\sqrt{(N_{\text{order}}+1)!}}\exp\left(\frac{(\sigma\Delta s)^{2}}{2}\right)$ (21)
The contraction factor $\sigma<1$ ensures convergence. ∎
Remark 1 (Comparison with Taylor Expansion).
The Taylor remainder $\mathbf{E}_{\text{truncation}}^{\text{Taylor}}=O((\Delta s)^{N+1}/(N+1)!)$ lacks the $\sqrt{(N+1)!}$ denominator and $\sigma^{N+1}$ suppression factor, resulting in looser bounds for non-monotonic trajectories.
A.3.2 Finite Difference Approximation Error
Lemma 2 (Finite Difference Approximation Error).
For a sufficiently smooth function $f\in C^{k+1}[0,T]$, the backward difference operator $\Delta^{(k)}$ satisfies:
$\|\Delta^{k}f(t)-f^{(k)}(t)\|=\frac{\Delta t_{\text{hist}}}{k+1}\|f^{(k+1)}(\xi)\|+O(\Delta t_{\text{hist}}^{2})$ (22)
for some $\xi\in[t-k\Delta t_{\text{hist}},t]$.
Lemma 3 (Gaussian Decay in Feature Dynamics).
Under the Gaussian feature dynamics (Observation 1), the expected norm of derivative approximations satisfies:
$\mathbb{E}[\|\Delta^{k}\mathbf{F}\|^{2}]=\frac{\sigma_{\mathbf{F}}^{2}}{\Gamma(k/2+1)}\cdot(1+O(k^{-1}))$ (23)
where $\sigma_{\mathbf{F}}^{2}$ is the feature variance and $\Gamma$ is the gamma function.
Lemma 4 (Total Approximation Error).
The cumulative finite difference error in HiCache satisfies:
$\mathbf{E}_{\text{approximation}}\leq C_{4}\Delta t_{\text{hist}}\sqrt{N_{\text{order}}}$ (24)
Proof.
Combining Lemma 2 and Lemma 3:
$\mathbf{E}_{\text{approximation}}$ $=\sum_{k=1}^{N_{\text{order}}}\mathbb{E}\left[\|\Delta^{k}f-f^{(k)}\|\cdot\frac{|\tilde{H}_{k}(\Delta s)|}{k!}\right]$ (25)
$\leq\sum_{k=1}^{N_{\text{order}}}\frac{C_{3}\Delta t_{\text{hist}}}{k+1}\cdot\frac{\sigma^{k}}{\sqrt{\Gamma(k/2+1)}}$ (26)
$\leq C_{4}\Delta t_{\text{hist}}\sum_{k=1}^{N_{\text{order}}}\frac{1}{k^{3/2}}$ (27)
$\leq C_{4}\Delta t_{\text{hist}}\sqrt{N_{\text{order}}}$ (28)
The rapid decay from both the Gaussian property and factorial normalization ensures convergence. ∎
A.3.3 Numerical Stability Analysis
The recurrence relation $H_{k+1}(x)=2xH_{k}(x)-2kH_{k-1}(x)$ becomes numerically unstable when $|x|\gg 1$. The contraction factor $\sigma$ limits the effective computation domain to $|\sigma\Delta s|<2$, maintaining condition number $\kappa\approx O(1)$.
Due to the validated Gaussian properties of feature differences, Hermite coefficients possess optimal decay properties:
$\text{Var}(c_{k})\propto\frac{1}{k!},\quad\text{Cov}(c_{i},c_{j})=0\text{ for }i\neq j$ (29)
Lemma 5 (Numerical Error Bound).
The floating-point error in HiCache is bounded by:
$\mathbf{E}_{\text{numerical}}\leq C_{5}\epsilon_{\text{machine}}\sqrt{\sum_{k=0}^{N_{\text{order}}}\frac{\sigma^{2k}}{k!}}\leq C_{5}\epsilon_{\text{machine}}e^{\sigma^{2}/2}$ (30)
A.3.4 Total Error Bound and Performance Advantages
Theorem 1 (Main Error Theorem - Complete Proof of Proposition 1 in the main paper).
Combining Lemma 1, Lemma 4, and Lemma 5, the total HiCache prediction error satisfies:
$\|\mathbf{E}_{\text{total}}\|$ $\leq\underbrace{C_{2}\frac{(\sigma\sqrt{2}|\Delta s|)^{N_{\text{order}}+1}}{\sqrt{(N_{\text{order}}+1)!}}e^{(\sigma\Delta s)^{2}/2}}_{\text{Truncation}}$
$\quad+\underbrace{C_{4}\Delta t_{\text{hist}}\sqrt{N_{\text{order}}}}_{\text{Approximation}}+\underbrace{C_{5}\epsilon_{\text{machine}}}_{\text{Numerical}}$ (31)
Corollary 3 (Comparative Advantage).
HiCache achieves superior error bounds compared to Taylor expansion through:
  1. 1.
    Faster convergence: $O(1/\sqrt{(N+1)!})$ vs. $O(1/(N+1)!)$
  2. 2.
    Exponential suppression: Factor $\sigma^{N+1}$ with $\sigma<1$
  3. 3.
    Adaptive basis: Hermite polynomials naturally match Gaussian feature statistics
Proof of Superiority.
The key advantage stems from the denominator $\sqrt{(N+1)!}$ in the Hermite truncation error, which grows much faster than $(N+1)!$ in the Taylor case. Combined with the $\sigma^{N+1}$ suppression factor, this yields exponentially tighter bounds for large $N$. ∎
A.4 Detailed Experimental Setup
This section provides comprehensive technical details for all experimental configurations mentioned in Section 4.1.
A.4.1 Model and Task Specifications
Text-to-Image Generation (FLUX.1-dev):
We employ the FLUX.1-dev model for high-resolution text-to-image synthesis. Images are generated at 1024×1024 resolution using 1,632 high-quality prompts sourced from the DrawBench benchmark. Quality assessment is performed using ImageReward, a robust perceptual metric for text-to-image alignment.
Text-to-Video Generation (Hunyuan-DiT):
The Hunyuan-DiT model is used for temporal content generation. We generate 4,730 video clips from 946 diverse prompts, evaluating performance using the comprehensive VBench framework which assesses multiple aspects including temporal consistency, object persistence, and text-video alignment.
Class-Conditional Image Generation (DiT-XL/2):
Following standard academic protocols, we conduct experiments on ImageNet using the DiT-XL/2 architecture. We generate 50,000 samples across all ImageNet classes and report FID-50k, sFID, and Inception Score for comprehensive quality assessment.
Image Super-Resolution (Inf-DiT):
We employ a modified patch-based Inf-DiT model for super-resolution tasks. Evaluation is conducted on the DIV8K dataset using standard PSNR and SSIM metrics to assess restoration quality and perceptual fidelity.
A.4.2 Hardware and Computational Resources
All experiments are conducted on enterprise-grade GPU infrastructure:
  • FLUX.1-dev experiments: NVIDIA H800 GPU
  • DiT-XL/2 experiments: NVIDIA A800 GPU
  • Hunyuan-DiT experiments: NVIDIA H100 GPU
  • Inf-DiT super-resolution: NVIDIA A100 GPU
A.4.3 Evaluation Protocols
Each task employs task-specific evaluation methodologies:
  • Text-to-Image: ImageReward scores for semantic alignment and image quality
  • Text-to-Video: VBench comprehensive evaluation including temporal consistency and motion quality
  • Class-Conditional: Standard ImageNet metrics (FID-50k, sFID, Inception Score)
  • Super-Resolution: Pixel-level metrics (PSNR, SSIM) for restoration accuracy
A.5 Super-Resolution Task: Detailed Experimental Results
This section provides comprehensive experimental results for HiCache’s evaluation on image super-resolution using the patch-based Inf-DiT model. The experiments were conducted using standard evaluation metrics including PSNR and SSIM for restoration quality assessment.
Table 5 presents detailed quantitative results across different acceleration intervals, comparing HiCache with TaylorSeer across multiple evaluation metrics including PSNR and SSIM.
Table 5: Performance Comparison of Different Guiders on Image Super-Resolution Task
Method Interval
Latency
Time
Speedup
FLOPs
FLOPs
Speedup
Restoration Track
[…]
HiCache
[…]
TaylorSeer
[…]
Note: HiCache uses $\sigma=0.5$. Values in (blue) denote the change relative to the interval=1 baseline.
The results demonstrate that HiCache maintains superior performance across all metrics, with particularly notable advantages at higher acceleration ratios. At interval=8, HiCache achieves a PSNR of 30.18 (only 0.69 drop from baseline), while TaylorSeer degrades to 29.71 (1.17 drop). This represents the first successful application of cache-based acceleration to super-resolution tasks, expanding the applicability of feature caching beyond traditional text-to-image and video generation.
A.6 Theoretical Analysis of Gaussian Feature Dynamics and Hermite Optimality
This section provides theoretical analysis supporting the core claims in Proposition 2 (HiCache Feature Prediction) in the main paper. We examine both the empirical foundation and theoretical justification for using Hermite polynomials as the prediction basis.
A.6.1 On the Gaussianity of Feature Derivatives (Observation 1 in the main paper)
While a complete theoretical proof of Gaussianity in neural network feature dynamics remains challenging, we provide several theoretical perspectives that support this empirical observation:
Central Limit Theorem Perspective
Consider the feature update through a transformer block:
$\mathbf{F}_{t+1}=\mathbf{F}_{t}+\text{MHSA}(\mathbf{F}_{t})+\text{MLP}(\mathbf{F}_{t})$ (32)
The finite difference approximations aggregate many such updates:
$\Delta^{k}\mathbf{F}_{t}=\frac{1}{\Delta t^{k}}\sum_{i=1}^{k}(-1)^{k-i}\binom{k}{i}\mathbf{F}_{t-i\Delta t}$ (33)
Proposition 4 (Asymptotic Gaussianity).
Under mild regularity conditions, if the individual feature updates are weakly dependent with finite variance, the normalized finite differences converge in distribution to a Gaussian:
$\frac{\Delta^{k}\mathbf{F}_{t}-\mathbb{E}[\Delta^{k}\mathbf{F}_{t}]}{\sqrt{\text{Var}[\Delta^{k}\mathbf{F}_{t}]}}\xrightarrow{d}\mathcal{N}(0,I)$ (34)
as the feature dimension $d\to\infty$.
Justification: This follows from the multivariate central limit theorem for weakly dependent random vectors. The transformer’s residual connections and normalization layers help maintain the required regularity conditions.
Random Matrix Theory Perspective
The attention mechanism involves random-like weight matrices due to the softmax operation on query-key similarities:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (35)
Lemma 6 (Gaussian Universality in Random Matrices).
For large transformer models with random initialization and sufficient depth, the distribution of feature differences approaches a universal form that, under appropriate scaling, exhibits Gaussian characteristics in the bulk of the spectrum.
This is related to the Wigner semicircle law and its generalizations, which predict Gaussian-like behavior in high-dimensional random matrix ensembles.
Diffusion Process Approximation
In the continuous-time limit, the discrete diffusion process can be approximated by a stochastic differential equation:
$d\mathbf{F}_{t}=\mu(\mathbf{F}_{t},t)dt+\sigma(\mathbf{F}_{t},t)dW_{t}$ (36)
where $W_{t}$ is a Wiener process. Under appropriate conditions on the drift $\mu$ and diffusion $\sigma$ coefficients, the marginal distributions of such processes exhibit Gaussian characteristics, especially for the increments.
A.6.2 Hermite Polynomials as Optimal Basis (Corollary 1 in the main paper)
Given the empirically validated Gaussian nature, we now establish why Hermite polynomials form an optimal basis:
Theorem 2 (Karhunen-Loève Expansion for Gaussian Processes).
Let $X(t)$ be a zero-mean Gaussian process with covariance kernel $K(s,t)$. If $K$ has the form:
$K(s,t)=\exp\left(-\frac{(s-t)^{2}}{2\tau^{2}}\right)$ (37)
then the eigenfunctions of the covariance operator are Hermite functions, and the optimal $L^{2}$ expansion is:
$X(t)=\sum_{n=0}^{\infty}\xi_{n}\psi_{n}(t)$ (38)
where $\psi_{n}$ are scaled Hermite functions and $\xi_{n}$ are uncorrelated Gaussian coefficients.
Proof Sketch.
The covariance operator $\mathcal{K}$ defined by:
$(\mathcal{K}f)(t)=\int K(s,t)f(s)ds$ (39)
has Hermite functions as eigenfunctions when $K$ has Gaussian form. This follows from the fact that Hermite functions are eigenfunctions of the Fourier transform composed with multiplication by a Gaussian, which is essentially what the Gaussian kernel operator represents. ∎
Corollary 4 (Optimality for Feature Prediction).
If feature differences $\Delta^{k}\mathbf{F}$ follow a multivariate Gaussian distribution, then the Hermite polynomial basis minimizes the expected squared prediction error:
$\mathbb{E}\left[\|\mathbf{F}_{t-\Delta t}-\sum_{k=0}^{N}c_{k}\phi_{k}(\Delta t)\|^{2}\right]$ (40)
among all orthogonal polynomial bases $\{\phi_{k}\}$.
Proof Sketch.
By the Karhunen-Loève theorem, the optimal basis for representing a stochastic process is given by the eigenfunctions of its covariance operator. For Gaussian processes with appropriate covariance structure, these are Hermite functions. The optimality follows from the fact that this expansion minimizes the mean squared error for any finite truncation order $N$. ∎
A.6.3 Robustness Under Approximate Gaussianity
Even if the Gaussianity assumption is only approximately satisfied, Hermite polynomials maintain advantages:
Proposition 5 (Robustness to Non-Gaussianity).
Let $p(\mathbf{x})$ be a distribution with finite moments that can be expressed as:
$p(\mathbf{x})=\phi(\mathbf{x})(1+\epsilon h(\mathbf{x}))$ (41)
where $\phi(\mathbf{x})$ is the Gaussian density, $|h(\mathbf{x})|\leq C$, and $\epsilon$ is small. Then the relative efficiency of Hermite basis compared to monomials degrades at most as $O(\epsilon)$.
This robustness property ensures that even under model misspecification, Hermite polynomials retain their advantages over standard polynomial bases.
A.6.4 Summary and Practical Implications
While we cannot provide a complete first-principles proof that neural network features must exhibit Gaussian dynamics, we have shown:
1. **Theoretical Support**: Multiple theoretical frameworks (CLT, random matrix theory, diffusion approximations) suggest Gaussianity is a reasonable expectation in high-dimensional neural networks.
2. **Conditional Optimality**: Given the empirically validated Gaussian nature, the optimality of Hermite polynomials follows rigorously from the Karhunen-Loève theorem.
3. **Robustness**: Even under approximate Gaussianity, Hermite bases maintain advantages over monomials due to their oscillatory nature and better numerical conditioning.
These theoretical insights, combined with our extensive empirical validation (Section 5.1), provide strong justification for the HiCache design choice of using Hermite polynomials as the prediction basis.
A.7 Implementation Details
This section provides the complete Python implementation of the HiCache algorithm, including all core functions for derivative approximation, feature prediction, and cache management.
1from typing import Dict, List, Union
2import torch
3import math
4
5def _hermite_polynomial(x: torch.Tensor, n: int) -> torch.Tensor:
6 """
7 Physicist’s Hermite polynomial H_n(x).
[…]
9 """
10 if n == 0:
11 return torch.ones_like(x)
12 elif n == 1:
13 return 2 * x
14
15 H_prev = torch.ones_like(x)
16 H_curr = 2 * x
17
18 for k in range(2, n + 1):
19 H_next = 2 * x * H_curr - 2 * (k - 1) * H_prev
20 H_prev, H_curr = H_curr, H_next
21
22 return H_curr
23
24def _collect_trajectory_feature(
25 cache_dic: Dict,
26 current: Dict,
27 feature: torch.Tensor
28):
29 """
30 Feature trajectory collector - automatically collects features during
31 cache updates. Supports simultaneous collection from multiple layers
32 and modules (collects from all required modules in a single inference
33 pass).
34
35 :param cache_dic: Cache dictionary
36 :param current: Information of the current step
37 :param feature: Feature tensor to collect
38 """
39 config = cache_dic.get(’feature_collection_config’, {})
40
41 # Support multi-layer feature collection
42 target_layers = config.get(’target_layers’, [14])
43 if isinstance(target_layers, int):
44 target_layers = [target_layers] # For backward compatibility
45
46 # Check collection conditions - collect only if in the target layer list
47 if current[’layer’] not in target_layers:
48 return
49
50 # NEW: Support simultaneous collection from multiple modules
51 target_modules = config.get(’target_modules’, [’any’])
52 if isinstance(target_modules, str):
53 target_modules = [target_modules]
54
55 # Check if the module needs to be collected
56 if ’any’ not in target_modules and current[’module’] not in target_modules:
57 return
58
59 # Support multi-stream collection
60 target_streams = config.get(’target_streams’, [’any’])
61 if isinstance(target_streams, str):
62 target_streams = [target_streams]
63
64 # Check if the stream needs to be collected
65 if ’any’ not in target_streams and current[’stream’] not in target_streams:
66 return
67
68 # NEW: Store by layer-module combination, supporting simultaneous
69 # collection from multiple modules
70 if ’trajectory_features’ not in cache_dic:
71 cache_dic[’trajectory_features’] = {}
72 cache_dic[’trajectory_metadata’] = {}
73
74 layer_key = current[’layer’]
75 module_key = current[’module’]
76
77 # Initialize layer-level storage
78 if layer_key not in cache_dic[’trajectory_features’]:
79 cache_dic[’trajectory_features’][layer_key] = {}
80 cache_dic[’trajectory_metadata’][layer_key] = {}
81
82 # Initialize module-level storage
83 if module_key not in cache_dic[’trajectory_features’][layer_key]:
84 cache_dic[’trajectory_features’][layer_key][module_key] = []
85 cache_dic[’trajectory_metadata’][layer_key][module_key] = []
86
87 # Collect features into the corresponding module storage
88 cache_dic[’trajectory_features’][layer_key][module_key].append(
89 feature.clone().detach().cpu()
90 )
91 cache_dic[’trajectory_metadata’][layer_key][module_key].append({
92 ’step’: current[’step’],
93 ’timestep’: current.get(’t’, 0),
94 ’cache_type’: current.get(’type’, ’full’),
95 ’layer’: current[’layer’],
96 ’module’: current[’module’],
97 ’stream’: current[’stream’]
98 })
99
100def derivative_approximation(
101 cache_dic: Dict,
102 current: Dict,
103 feature: torch.Tensor
104):
105 """
106 Compute derivative approximation.
107
108 :param cache_dic: Cache dictionary
109 :param current: Information of the current step
110 """
111 # NEW: Feature collection hook
112 if cache_dic.get(’enable_feature_collection’, False):
113 _collect_trajectory_feature(cache_dic, current, feature)
114 # If only for feature collection, skip the rest of the cache operations
115 if not cache_dic.get(’taylor_cache’, False):
116 return
117
118 # Safety check: ensure cache structure exists
119 try:
120 cache_path = cache_dic[’cache’][-1][current[’stream’]][current[’layer’]]
121 cache_module = cache_path[current[’module’]]
122 except KeyError:
123 return
124
125 # Safety check: ensure there are enough activated_steps to calculate the difference
126 if len(current[’activated_steps’]) < 2:
127 # If there is only one or no activated_steps, the difference cannot be calculated, so return directly
128 return
129
130 activated_steps = current[’activated_steps’]
131 difference_distance = activated_steps[-1] - activated_steps[-2]
132 #activated_times = current[’activated_times’]
133 #difference_distance = activated_times[-1] - activated_times[-2]
134
135 updated_taylor_factors = {}
136 updated_taylor_factors[0] = feature
137
138 for i in range(cache_dic[’max_order’]):
139 has_cache = cache_module.get(i, None) is not None
140 step_condition = current[’step’] > cache_dic[’first_enhance’] - 2
141 if has_cache and step_condition:
142 numerator = updated_taylor_factors[i] - cache_module[i]
143 updated_taylor_factors[i + 1] = numerator / difference_distance
144 else:
145 break
146
147 cache_path[current[’module’]] = updated_taylor_factors
148
149def get_collected_features(cache_dic: Dict) -> tuple:
150 """
151 Get the collected feature trajectories - supports multi-layer data.
152
153 :param cache_dic: Cache dictionary
154 :return: (features_dict, metadata_dict) tuple where keys are layer indices
155 """
156 features = cache_dic.get(’trajectory_features’, {})
157 metadata = cache_dic.get(’trajectory_metadata’, {})
158 return features, metadata
159
160def clear_collected_features(cache_dic: Dict):
161 """
162 Clear the collected feature trajectories.
163
164 :param cache_dic: Cache dictionary
165 """
166 if ’trajectory_features’ in cache_dic:
167 del cache_dic[’trajectory_features’]
168 if ’trajectory_metadata’ in cache_dic:
169 del cache_dic[’trajectory_metadata’]
170
171
172def taylor_formula(cache_dic: Dict, current: Dict) -> torch.Tensor:
173 """
174 Feature prediction dispatcher: chooses between Taylor or Hermite prediction.
175
176 :param cache_dic: Cache dictionary
177 - ’prediction_mode’: ’taylor’ or ’hermite’. Defaults to ’taylor’ if not specified.
178 - ’use_hermite’: (Legacy) If True and ’prediction_mode’ is not set, mode becomes ’hermite’.
179 :param current: Information of the current step
180 """
181 # Determine prediction mode, with backward compatibility for ’use_hermite’
182 if ’prediction_mode’ in cache_dic:
183 mode = cache_dic[’prediction_mode’]
184 elif cache_dic.get(’use_hermite’, False):
185 mode = ’hermite’
186 else:
187 mode = ’taylor’
188
189 # Dispatch based on mode
190 if mode == ’taylor’:
191 return _taylor_expansion_formula(cache_dic, current)
192
193 elif mode == ’hermite’:
194 return _hermite_prediction_formula(cache_dic, current)
195
196 else:
197 error_msg = (f"Unknown prediction_mode: ’{mode}’. "
198 f"Must be ’taylor’ or ’hermite’.")
199 raise ValueError(error_msg)
200
201
202def _taylor_expansion_formula(cache_dic: Dict, current: Dict) -> torch.Tensor:
203 """
204 Standard Taylor expansion prediction.
205 Uses a power function basis: F_pred = F_0 + Sum (1/k!) * x^k * Delta^kF
206 """
207 activated_steps = current[’activated_steps’]
208 x = current[’step’] - activated_steps[-1]
209 #x = current[’t’] - current[’activated_times’][-1]
210 output = 0
211
212 # FIX: Safety check, ensure cache structure exists
213 try:
214 cache_root = cache_dic[’cache’][-1]
215 stream_cache = cache_root[current[’stream’]]
216 layer_cache = stream_cache[current[’layer’]]
217 feats_d = layer_cache[current[’module’]]
218 except KeyError:
219 # FIX: If the cache does not exist, it means this is the first access to this module.
220 # In this case, we should initialize the cache structure.
221 cache_root = cache_dic[’cache’][-1]
222 if current[’stream’] not in cache_root:
223 cache_root[current[’stream’]] = {}
224 stream_cache = cache_root[current[’stream’]]
225 if current[’layer’] not in stream_cache:
226 stream_cache[current[’layer’]] = {}
227 layer_cache = stream_cache[current[’layer’]]
228 if current[’module’] not in layer_cache:
229 layer_cache[current[’module’]] = {}
230
231 # FIX: If the cache is empty, it means this is the first call to this module.
232 # In this case, we should throw a more user-friendly error message,
233 # guiding the user to check the configuration.
234 error_msg = (f"Cache not found for stream=’{current[’stream’]}’, "
235 f"layer={current[’layer’]}, module=’{current[’module’]}’. "
236 f"This usually means the first step was not run in ’full’ "
237 f"mode to initialize the cache. Please check your cache "
238 f"configuration and ensure first_enhance >= 1.")
239 raise ValueError(error_msg)
240
241 # FIX: Use the max_order parameter to limit the number of terms used
242 max_order = cache_dic.get(’max_order’, 3)
243 # +1 because we include 0th order
244 effective_order = min(max_order + 1, len(feats_d))
245
246 for i in range(effective_order):
247 output += (1 / math.factorial(i)) * feats_d[i] * (x ** i)
248
249 return output
250
251
252def _hermite_prediction_formula(
253 cache_dic: Dict,
254 current: Dict
255) -> torch.Tensor:
256 """
257 Feature prediction based on Hermite polynomials.
258
259 Uses Hermite polynomials as basis functions instead of standard power
[…]
261
262 Where:
263 - H_k(x) is the k-th order Hermite polynomial
264 - Delta^kF is the k-th order difference feature
265 - x is the time step difference
266
267 Advantages of this method:
268 1. Hermite polynomials are orthogonal, providing better numerical
269 stability.
270 2. Higher approximation accuracy for certain types of functions.
271 3. The growth of the polynomial can be controlled by a scaling factor.
272 """
273 # Get the actual time step difference, consistent with the original
274 # Taylor expansion
275 activated_steps = current[’activated_steps’]
276 x = current[’step’] - activated_steps[-1]
277 # If time difference is needed, uncomment the line below
278 # x = current[’t’] - current[’activated_times’][-1]
279
280 # Get feature cache
281 try:
282 cache_root = cache_dic[’cache’][-1]
283 stream_cache = cache_root[current[’stream’]]
284 layer_cache = stream_cache[current[’layer’]]
285 feats_d = layer_cache[current[’module’]]
286 except KeyError:
287 error_msg = (f"Cache not found for stream=’{current[’stream’]}’, "
288 f"layer={current[’layer’]}, module=’{current[’module’]}’")
289 raise ValueError(error_msg)
290
291
292 # FIX: Use the max_order parameter to limit the order
293 max_order = cache_dic.get(’max_order’, 3)
294 # available order = number of historical terms - 1
295 available_order = len(feats_d) - 1
296 # use the smaller value
297 order = min(max_order, available_order)
298
299 if order < 1:
300 # Not enough history, return the latest feature
301 return feats_d.get(0)
302
303 # F_0
304 F_latest = feats_d[0].clone()
305
306 # Convert the time step difference to a tensor, keeping the same
307 # dtype and device as the feature
308 x_tensor = torch.tensor(
309 float(x),
310 dtype=F_latest.dtype,
311 device=F_latest.device
312 )
313
314 # Get the scaling factor to control the growth of the Hermite polynomial
315 scale_factor = cache_dic.get(’hermite_scale_factor’, 0.5)
316 x_scaled = x_tensor * scale_factor
317
318 # Construct Hermite prediction
319 pred = F_latest.clone()
320
321 for k in range(1, order + 1):
322 diff_k = feats_d[k]
323 Hk = _hermite_polynomial(x_scaled, k)
324 # Consider the effect of the scaling factor
325 alpha = float(Hk / math.factorial(k)) * (scale_factor ** k)
326 pred.add_(diff_k, alpha=alpha)
327
328 return pred
329
330
331
332def taylor_cache_init(cache_dic: Dict, current: Dict):
333 """
334 Initialize Taylor cache and allocate storage for different-order
335 derivatives in the Taylor cache.
336
337 :param cache_dic: Cache dictionary
338 :param current: Information of the current step
339 """
340 if (current[’step’] == 0) and (cache_dic[’taylor_cache’]):
341 cache_root = cache_dic[’cache’][-1]
342 stream_cache = cache_root[current[’stream’]]
343 layer_cache = stream_cache[current[’layer’]]
344 layer_cache[current[’module’]] = {}
Listing 1: Core Implementation of HiCache Algorithm
The implementation includes several key components:
  • Hermite Polynomial Computation (_hermite_polynomial): Recursive computation using the recurrence relation in Eq. (14)
  • Feature Collection (_collect_trajectory_feature): Multi-layer, multi-module feature trajectory collection
  • Derivative Approximation (derivative_approximation): Finite difference approximations as described in Algorithm 1
  • Prediction Dispatcher (taylor_formula): Interface supporting both Taylor and Hermite prediction modes
  • Hermite Prediction (_hermite_prediction_formula): Hermite polynomial-based feature prediction as described in Algorithm 2
  • Taylor Prediction (_taylor_expansion_formula): Taylor expansion implementation